Ishaan is 2 times as old as Christopher. 35 years ago, Ishaan was 7 times as old as Christopher.
How old is Ishaan now?

Found 4 solutions by Cromlix, stanbon, Alan3354, Vixiachase:
Answer by Cromlix(4381)

Hi there,
Ishaan is 2 times as old as Christopher.
I = 2C.......(1)
35 years ago, Ishaan was 7 times as old as Christopher.
I - 35 = 7(C - 35).......(2)
I - 35 = 7C - 245........(2)
Substitute 2C for I from Eq(1) in Eq(2)
2C - 35 = 7C - 245
Collect like terms:
2C - 7C = -245 + 35
-5C = -210
C = 42
Ishaan is 84 years old
Christopher is 42 years old
Hope this helps :-)

Answer by stanbon(75887)

Ishaan is 2 times as old as Christopher. 35 years ago, Ishaan was 7 times as old as Christopher.
How old is Ishaan now?
-------------
I = 2*C
I-35 = 7(C-35)
-------
Substitute for "I" and solve for "C"::
2C - 35 = 7C - 245
5C = 210
C = 42 yrs (Christopher's age now)
Ishanna's age now = 2C = 84 yrs
-----------
Cheers
Stan H.
-------------

Answer by Alan3354(69443)

Ishaan is 2 times as old as Christopher. 35 years ago, Ishaan was 7 times as old as Christopher.
How old is Ishaan now?
===========
I = 2C
I - 35 = 7(C-35)
Sub for I
-----
2C - 35 = 7C - 245
5C = 210
C = 42

Answer by Vixiachase(22)